7. A company receives $7 million from New Enterprise Associates (NEA) for a post-money evaluation of $20 million.

a. What is the Pre-money Valuation and what percentage of the company will be owned by NEA?

b. How can they calculate their ROI?
Answer #1

a)

Pre-money Valuation = Post Money Valuation-Investment by NEA = 20-7 = $13 Million

Percentage owned by NEA = Investment by NEA/Post Money Valuation = 7/20 = 35%

b)

NEA can calculate their ROI as follows:

Whenever the company is taken over, the Value at which it is taken over, will be considered.

ROI will be = [(New Value*NEA's share)-Investment]/Investment

Say for example, Company is taken over for $30 Million.]

In that case, ROI for NEA will be = [(30*0.35)-7]/7 = (10.5-7)/7 = 50%